OpenStudy (ivitt):

What is the next term in the pattern? 2, 6, 18, 54, 162, HALP!!

OpenStudy (ivitt):

Do I multiply by 3 or??

OpenStudy (mathmale):

Evaluate the following: 3*2, 3*6, 3*18, 3*54. What next?

OpenStudy (ivitt):

3*162 3*486??

OpenStudy (mathmale):

Good. What is your conclusion?

OpenStudy (ivitt):

Multiply by 3

OpenStudy (mathmale):

Need a bit more info than that. Given the nth term of a sequence, how does one find the (n+1)st term? \[a _{n+1}=3(a _{n})\]

OpenStudy (mathmale):

This says: "Given the term a sub n of a certain arithmetic sequence, find the next term by multiplying a_sub_n by 3."

OpenStudy (phi):

yes, the next term is 3*162 = 486 when I do these problems I look for a "constant multiple" (by dividing the second number by the number in front of it: example 6/2 =3 then try 18/6 = 3 if that does not work, I look for a constant difference (which is not the case here)